OpenStudy (anonymous):

y varies jointly with p and q. When p = 7 and q = 2, y = 70. Find y when p = 8 and q = 4

OpenStudy (amistre64):

stop deleting your questions, its rude to those who are reading it and helping out.

OpenStudy (anonymous):

true!!

OpenStudy (amistre64):

to vary jointly is code for multiplications: y = pq is the setup for this

OpenStudy (amistre64):

in order for the numbers to make any sense in this; we have to introduce a "constant" of variation that will control the situation

OpenStudy (anonymous):

y = 160

OpenStudy (amistre64):

y = kpq ; or rather k = y/(pq) will be needed for it to work out

OpenStudy (amistre64):

When p = 7 and q = 2, y = 70. k = 70/(7*2) = 70/14 = 35/7 = 5 Find y when p = 8 and q = 4 y = 5pq, y = 5(8*4) = 5(32) = 160 if i see it right

OpenStudy (anonymous):

y=160